Two electrons (-1.6 x 10^23 C) are separated by an insulator in a capacitor at a distance of 0.0001 m. What is the electronic force between these electrons? (k- 9.0 x 10^2 N m^2/ C)
1
Expert's answer
2011-01-31T12:38:01-0500
The Coulomb force of repulsion of two elestrons is
F = k q1 q2 / R2 = 9.0 x 109 * (-1.6 x 10-19)2 / 0.00012 = 2.30400 × 10-20 N
With k = 9.0 x 102 N m2/ C:
F = k q1 q2 / R2 = 9.0 x 102 * (-1.6 x 10-19)2 / 0.00012 = 2.30400 × 10-27 N